Velocity (mM/min)
[S], mM Uninhibited Inhibited
1.75 1.94 1.38
2.17 2.26 1.67
3.00 2.85 2.13
5.50 3.55 2.97
10.5 4.39 3.83

(a) Compute for the values of 1/[S] and 1/v for both the uninhibited and inhibited reactions.

1/Velocity (min/mM)
1/[S], mM-1 Uninhibited Inhibited

(b) Use MS Excel program to construct the Lineweaver-Burke plots for both reaction systems.Conduct linear regression analysis and compute for the equation of the lines and their respective R2 values. Compute for the Vmax and Km values of the uninhibited and inhibited reaction. Determine the mode of inhibition based on the generated graph and the values of the kinetic parameters.
Answer #1

(a):

(b):

Uninhibited Reaction:

Y = 0.596X + 0.167

In Lineweaver Burk plot (Uninhibited):

Y-intercept = 0.167 = 1/Vmax

=> Vmax = 1/0.167 = 5.99 mM/min (Answer)

X-intercept is when Y = 0

=> 0 = 0.596X + 0.167

=> X = - 0.167 / 0.596 = - 0.28020

Also X-intercept = - 1/Km

=> - 0.28020 = - 1/Km

=> Km = 1/0.28020 = 3.57 mM (Answer)

Inhibited Reaction:

Y = 0.965X + 0.161

In Lineweaver Burk plot (Inhibited):

Y-intercept = 0.161 = 1/Vmax

=> Vmax = 1/0.161 = 6.21 mM/min (Answer)

X-intercept is when Y = 0

=> 0 = 0.965X + 0.161

=> X = - 0.161 / 0.965 = - 0.16684

Also X-intercept = - 1/Km

=> - 0.16684 = - 1/Km

=> Km = 1/0.16684 = 5.99 mM (Answer)

Since Km is increased from uninhibited to inhibited, where as Vmax values are nearly same, this is a competitive inhibition.